# On fixed points of pseudo-Anosov maps

### v1

@@ 8.2. Proving shuffling
Suppose that $Y\subset S$ is qq–compatible and that $\mathring{W}\in\mathcal{F}_{Y}$ satisfies $d_{\mathring{W}}(\lambda^{-},\lambda^{+})\geq 4$. Then for each $\alpha$ in $\mathcal{AC}(Y)$, $\pi^{q}_{\mathring{W}}(\alpha)\neq\emptyset$.
@@
It remains to consider the situation where $Y=S$ is the whole surface. In this setting, we will once again base our proof on Lemma 8.12 and Lemma 8.15. The challenge is that $d_{S}(\lambda^{+},\lambda^{-})$ is infinite, so we must instead replace this distance by the stable translation length $\ell_{S}(f)$, defined via (2.2).
@@ 9. Examples and constructions
Moreover, if $\alpha$ is a component of the intersection $S_{1}\cap S_{2}$ that is essential in each $S_{i}$, then the smoothing $S_{3}$ contains two disjoint preimages of $\alpha$, which are permuted under the first return map $f_{3}$. In other words, there is a curve on $S_{3}$ which is disjoint from its image under $f_{3}$, hence $f_{3}$ is fixed-point-free and not strongly irreducible.
@@
Let $f\colon S\to S$ be a pseudo-Anosov map such that $Y\subset S$ is a non-annular, non-overlapped essential subsurface. That is, $f(Y)\cap Y=\emptyset$. In particular, since $Y$ is non-annular, $Y$ itself supports a pseudo-Anosov homeomorphism $g\colon Y\to Y$. Then just as above, for all but finitely many $n$, the map $f\circ g^{n}$ is isotopic to a pseudo-Anosov homeomorphism whose number of fixed points is no greater than that of $f$. Moreover, it is easy to see (e.g. using the work of Brock [9]) that the volume of the mapping tori of $f\circ g^{n}$ go to $\infty$.

### v2

@@ 8.2. Proving shuffling
Suppose that $Y\subset S$ is qq–compatible and that $\mathring{W}\in\mathcal{F}_{Y}$ satisfies $d_{\mathring{W}}(\lambda^{-},\lambda^{+})\geq 4$. Then for each $\alpha$ in $\mathcal{AC}(Y)$, we have $\pi^{q}_{\mathring{W}}(\alpha)\neq\emptyset$.
@@
It remains to consider the situation where $Y=S$ is the whole surface. In this setting, we will once again base our proof on Lemmas 8.12 and 8.15. The challenge is that $d_{S}(\lambda^{+},\lambda^{-})$ is infinite, so we must instead replace this distance by the stable translation length $\ell_{S}(f)$, defined via (2.2).
@@ 9. Examples and constructions
Moreover, if $\alpha$ is a component of the intersection $S_{1}\cap S_{2}$ that is essential in each $S_{i}$, then the smoothing $S_{3}$ contains two disjoint preimages of $\alpha$ – say, $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ – such that the first return map $f_{3}$ sends $\alpha^{\prime}$ to $\alpha^{\prime\prime}$. We conclude that $f_{3}$ is both fixed-point-free and not strongly irreducible.
@@
Let $f\colon S\to S$ be a pseudo-Anosov map such that $Y\subset S$ is a non-annular, non-overlapped essential subsurface. That is, $f(Y)\cap Y=\emptyset$. In particular, since $Y$ is non-annular, the interior of $Y$ itself supports a pseudo-Anosov homeomorphism $g$. Then just as above, for all but finitely many $n$, the map $f\circ g^{n}$ is isotopic to a pseudo-Anosov homeomorphism whose number of fixed points is no greater than that of $f$. Moreover, it follows from the work of Brock [9] that the volume of the mapping tori of $f\circ g^{n}$ go to $\infty$.